# Differential Privacy Analysis of Decentralized Gossip Averaging under Varying Threat Models

### v2

@@ Appendix E Proof Theorem 13
First, from the compact SVD we see that for any matrix $A$, the product $A^{+}A$ is the orthogonal projector onto $\mathrm{row}(A)$. Set $P_{T}:=H_{T}^{+}H_{T}$ to be the orthogonal projection onto row space of $H_{T}$. Let $c\in\{\pm 1\}^{T}$ and define $z:=c\otimes e_{j}\in\mathbb{R}^{nT}$. Our strategy is to study the sensitivity
$\Delta_{T}(c,j\to i)=\|P_{T}z\|_{2}$
and derive a bound for it that is independent of $c$.
Because $W$ is primitive and row-stochastic, by Lemma 11 there exist constants $M>0$ and $\rho\in(0,1)$ such that
$W^{t}=\Pi+R^{t},\qquad\Pi:=\mathbf{1}\pi^{\top},\qquad\|R^{t}\|_{2}\leq M\rho^{t}\quad\forall t\geq 0.$ (E.1)
The $(t,s)$ block of $H_{T}$ (with $0\leq s\leq t\leq T-1$) is given by
$H_{T}(t,s)=e_{i}^{\top}W^{t-s}=e_{i}^{\top}\Pi+e_{i}^{\top}R^{t-s}=\pi^{\top}+e_{i}^{\top}R^{t-s}.$
Define the stationary part of $H_{T}$ as $\bar{H}_{T}$ such that blocks of $\bar{H}_{T}$ are given as
$\bar{H}_{T}(t,s)=\pi^{\top}\quad\text{ for }s\leq t,\qquad\bar{H}_{T}(t,s)=0\quad\text{for }s>t.$
Furthermore, define the transient part
$K_{T}:=H_{T}-\bar{H}_{T}.$
Thus
$H_{T}=\bar{H}_{T}+K_{T}.$
We clearly have that
$\mathrm{row}(\bar{H}_{T})=\{(\alpha_{0}\pi,\alpha_{1}\pi,\dots,\alpha_{T-1}\pi)\in\mathbb{R}^{nT}:\alpha_{t}\in\mathbb{R}\}$
and we can take as an orthonormal basis for $\mathrm{row}(\bar{H}_{T})$ for example $\{e_{1}\otimes\pi/\|\pi\|_{2},\ldots,e_{T}\otimes\pi/\|\pi\|_{2}\}.$ Therefore the orthogonal projection onto $\mathrm{row}(\bar{H}_{T})$ is given by
$\bar{P}_{T}z=\sum_{\ell=1}^{T}\langle e_{\ell}\otimes\pi/\|\pi\|_{2},c\otimes e_{j}\rangle=c\;\otimes\;\frac{\pi_{j}}{\|\pi\|_{2}^{2}}\pi.$
and we have that
$\|\bar{P}_{T}z\|_{2}^{2}=\sum_{t=0}^{T-1}\left\|c_{t}\,\frac{\pi_{j}}{\|\pi\|_{2}^{2}}\pi\right\|_{2}^{2}=T\,\frac{\pi_{j}^{2}}{\|\pi\|_{2}^{2}}.$
We next study the transient part and show using standard perturbation theory for projection matrices that the distance between the projectors $P_{T}$ and $\bar{P}_{T}$ satisfies
$\|P_{T}-\bar{P}_{T}\|_{2}=O(1/\sqrt{T}).$
This rate is derived using the Wedin theorem which states that the distance between the row spaces $\mathrm{row}(H_{T})$ and $\mathrm{row}(\bar{H}_{T})$ is a function of $\|K_{T}\|_{2}$ (since the perturbation $H_{T}-\bar{H}_{T}=K_{T}$) and the smallest non-zero singular value of $\bar{H}_{T}$, $\sigma_{\min}(\bar{H}_{T})$.
From the Toeplitz structure it follows that clearly $\|K_{T}\|_{2}$ equals the 2-norm of the first column block of $K_{T}$. Moreover, from the transient behavior stated in Eq. (E.1) and a geometric sum it follows that $\|K_{T}\|_{2}=O(1)$. We see that
$\bar{H}_{T}=U\otimes\pi^{\top}$
for
$U=\begin{pmatrix}1&0&\cdots&0\\ 1&1&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 1&1&\cdots&1\end{pmatrix}$.
The singular values of a Kronecker product $A\otimes B$ are the products of the singular values of $A$ and $B$. Thus, the non-zero singular values of $\bar{H}_{T}$ are:
$\sigma_{k}(\bar{H}_{T})=\sigma_{m}(U)\cdot\sigma_{l}(\pi^{\top}).$
Since $\pi^{\top}$ has rank $1$, $\sigma_{1}(\pi^{\top})=\|\pi^{\top}\|_{2}=\|\pi\|_{2}.$ We also know that the smallest singular value of $U$ is given by
$\sigma_{\min}(U)=\sqrt{T}.$
Thus, $\sigma_{\min}(\bar{H}_{T})=\sqrt{T}\|\pi\|_{2}$, and by the Wedin theorem we have that
$\|P_{T}-\bar{P}_{T}\|_{2}\leq\frac{2\|K_{T}\|_{2}}{\sigma_{\min}(\bar{H}_{T})}=O(1/\sqrt{T}).$
Let $\mathcal{E}_{T}=P_{T}-\bar{P}_{T}$, so $\|\mathcal{E}_{T}\|_{2}=O(1/\sqrt{T})$.
We expand $\|P_{T}z\|_{2}^{2}$ using $P_{T}=\bar{P}_{T}+\mathcal{E}_{T}$. Recall $z=c\otimes e_{j}$ and $\|z\|_{2}=\sqrt{T}$:
$\|P_{T}z\|_{2}^{2}=\|(\bar{P}_{T}+\mathcal{E}_{T})z\|_{2}^{2}=\|\bar{P}_{T}z\|_{2}^{2}+2(\bar{P}_{T}z)^{\top}(\mathcal{E}_{T}z)+\|\mathcal{E}_{T}z\|_{2}^{2}.$ (E.2)
Using a norm bound on $\mathcal{E}_{T}$, we have
$|2(\bar{P}_{T}z)^{\top}(\mathcal{E}_{T}z)|$ $\leq 2\|\bar{P}_{T}z\|_{2}\|\mathcal{E}_{T}z\|_{2}$
$\leq 2\|\bar{P}_{T}z\|_{2}\|\mathcal{E}_{T}\|_{2}\|z\|_{2}$
$\leq 2\left(\sqrt{T}\frac{\pi_{j}}{\|\pi\|_{2}}\right)\cdot O(1/\sqrt{T})\cdot\sqrt{T}=O(\sqrt{T}).$
Furthermore,
$\|\mathcal{E}_{T}z\|_{2}^{2}\leq\|\mathcal{E}_{T}\|_{2}^{2}\|z\|_{2}^{2}=O(1/T)\cdot T=O(1).$
[…]

### v3

@@ G.2 Bound for the Transient Term
For the proof we repeatedly use the Schur complement (Gallier, 2010). For $t=1,\dots,T$, let $K_{t}$ be the leading $t\times t$ principal submatrix of $K_{T}$ and write
$K_{t}=\begin{pmatrix}K_{t-1}&k_{t}\\ k_{t}^{\top}&\kappa_{t}\end{pmatrix},\qquad s_{t}:=\kappa_{t}-k_{t}^{\top}K_{t-1}^{-1}k_{t}>0.$
For any $x\in\mathbb{R}^{T}$ and $t\in[T]$, define the residual
$r_{t}(x):=x_{t}-k_{t}^{\top}K_{t-1}^{-1}x_{1:t-1}.$
A standard block inverse identity with Schur complement gives
$x^{\top}K_{T}^{-1}x=\sum_{t=1}^{T}\frac{r_{t}(x)^{2}}{s_{t}}.$ (G.15)
We next lower bound $s_{t}$ uniformly. Let $h_{t}^{\top}$ be the $t$th row of $H_{T}$. Then $K_{T}$ is the Gram matrix of $\{h_{t}\}_{t=1}^{T}$, and $s_{t}$ equals the squared distance from $h_{t}$ to $\mathrm{span}\{h_{1},\dots,h_{t-1}\}$. Since $H_{T}$ is block lower-triangular with diagonal block $S=e_{i}^{\top}$, each diagonal component is orthogonal to $\mathrm{span}\{h_{1},\dots,h_{t-1}\}$ and we clearly have that
$s_{t}\geq\|S\|_{2}^{2}=1.$ (G.16)
Combining (G.15)–(G.16) gives
$x^{\top}K_{T}^{-1}x\leq\sum_{t=1}^{T}r_{t}(x)^{2}.$ (G.17)
We now bound the residuals for $x=b^{(R)}(c)$. Since $k_{t}^{\top}K_{t-1}^{-1}$ is the minimizer of $\alpha^{\top}\mapsto(x_{t}-\alpha^{\top}x_{1:t-1})^{2}$, for any $\widehat{x}_{t}=\alpha^{\top}x_{1:t-1}$,
$r_{t}(x)^{2}\leq(x_{t}-\widehat{x}_{t})^{2}.$ (G.18)
Fix $t\geq 2$ and choose a truncation length
$m_{t}:=\left\lceil\frac{\log t}{\log(1/\rho)}\right\rceil\quad\Rightarrow\quad\rho^{m_{t}}\leq\frac{1}{t}.$
Define a predictor $\widehat{b}^{(R)}_{t}$ as
$\widehat{b}^{(R)}_{t}:=\sum_{q=1}^{m_{t}}\beta_{q}\,b^{(R)}_{t-q}(c),\qquad\beta_{q}:=(W^{q})_{ii}.$ (G.19)
We show that the error $e_{t}:=b^{(R)}_{t}-\widehat{b}^{(R)}_{t}$ is $O\!\big(\rho^{m_{t}}/(1-\rho)\big)$. Using (G.10) we may write
$b^{(R)}_{t}(c)=\sum_{m=1}^{t-1}r_{m}\,c_{t-m},\qquad r_{m}:=(R^{m})_{ij}.$
Similarly,
$b^{(R)}_{t-q}(c)=\sum_{m=1}^{t-q-1}r_{m}\,c_{t-q-m}=\sum_{\ell=q+1}^{t-1}r_{\ell-q}\,c_{t-\ell},$
and therefore
$\widehat{b}^{(R)}_{t}=\sum_{q=1}^{m_{t}}\beta_{q}\sum_{\ell=q+1}^{t-1}r_{\ell-q}\,c_{t-\ell}=\sum_{\ell=2}^{t-1}\left(\sum_{q=1}^{\min\{m_{t},\ell-1\}}\beta_{q}r_{\ell-q}\right)c_{t-\ell}.$
Hence
$e_{t}=\sum_{\ell=1}^{t-1}\left(r_{\ell}-\sum_{q=1}^{\min\{m_{t},\ell-1\}}\beta_{q}r_{\ell-q}\right)c_{t-\ell}.$ (G.20)
Now use the identity $W^{q}R^{\ell-q}=R^{\ell}$ for all $1\leq q\leq\ell-1$, which follows from $W=P+R$ and $PR=RP=0$. Taking the $(i,j)$ entry,
$r_{\ell}=(R^{\ell})_{ij}=\sum_{u=1}^{n}(W^{q})_{iu}\,(R^{\ell-q})_{uj}.$ (G.21)
Recalling $\beta_{q}=(W^{q})_{ii}$ gives
$r_{\ell}-\beta_{q}r_{\ell-q}=\sum_{u\neq i}(W^{q})_{iu}\,(R^{\ell-q})_{uj}.$ (G.22)
Summing (G.22) over $q=1,\dots,\min\{m_{t},\ell-1\}$ and plugging into (G.20), we see that for all $\ell\geq m_{t}+1$ the coefficient of $c_{t-\ell}$ is a sum of terms $(W^{q})_{iu}(R^{\ell-q})_{uj}$ with $u\neq i$ and $1\leq q\leq m_{t}$. Using row-stochasticity ($\sum_{u\neq i}(W^{q})_{iu}\leq 1$) and $|(R^{k})_{uj}|\leq\|R^{k}\|_{2}\leq\rho^{k}$, we obtain
$\left|r_{\ell}-\sum_{q=1}^{m_{t}}\beta_{q}r_{\ell-q}\right|\leq\sum_{q=1}^{m_{t}}\sum_{u\neq i}(W^{q})_{iu}\,|(R^{\ell-q})_{uj}|\leq\sum_{q=1}^{m_{t}}\rho^{\ell-q}\leq\frac{\rho^{\ell-m_{t}}}{1-\rho}.$
For $\ell\leq m_{t}$ the bracket in (G.20) is uniformly bounded by
$|r_{\ell}|+\sum_{q\leq\ell-1}|\beta_{q}||r_{\ell-q}|\leq\rho^{\ell}+\sum_{q\leq\ell-1}\rho^{\ell-q}\leq 1/(1-\rho).$
Combining these bounds and using $|c_{t-\ell}|=1$ gives
$|e_{t}|\leq\sum_{\ell=m_{t}+1}^{t-1}\frac{\rho^{\ell-m_{t}}}{1-\rho}\leq\frac{1}{(1-\rho)^{2}}\rho^{m_{t}}\leq\frac{1}{(1-\rho)^{2}}\frac{1}{t}.$ (G.23)
Finally, by (G.18) and (G.23),
$r_{t}(b^{(R)}(c))^{2}\leq e_{t}^{2}\leq\frac{1}{(1-\rho)^{4}}\frac{1}{t^{2}}.$ (G.24)
Using (G.17), summing over $t$ and using the bound (G.24) gives
$b^{(R)}(c)^{\top}K_{T}^{-1}b^{(R)}(c)\leq\sum_{t=1}^{T}r_{t}(b^{(R)}(c))^{2}\leq\frac{1}{(1-\rho)^{4}}\sum_{t=1}^{\infty}\frac{1}{t^{2}}\leq\frac{\pi^{2}/6}{(1-\rho)^{4}}.$
This proves the claim. ∎
Appendix H Extension to Multiple Observed Nodes
We next give proof of Theorem 15. The proof follows exactly the same structure as the proof of Theorem 13: we express the sensitivity as a quadratic form, decompose the dynamics into a stationary component and a transient component, show that the stationary part gives the term linear in $T$, and prove via an auxiliary technical lemma that transient contributions are uniformly bounded in $T$.
H.1 Proof of Theorem 15
Proof.
We start by stating the sensitivity formula as a quadratic form. For $c\in\{\pm 1\}^{T}$ define
$v(c):=c\otimes e_{j},\qquad b(c):=H_{T}v(c)\in\mathbb{R}^{mT},$
and set
$K_{T}:=H_{T}H_{T}^{\top}.$
Since $H_{T}^{+}=H_{T}^{\top}(H_{T}H_{T}^{\top})^{-1}$ and $H_{T}$ has full row rank, $K_{T}\succ 0$ and
$\|H_{T}^{+}H_{T}v(c)\|_{2}^{2}=b(c)^{\top}K_{T}^{-1}b(c),$
hence
$\Delta_{T}(j\to\mathcal{A})^{2}=\max_{c\in\{\pm 1\}^{T}}b(c)^{\top}K_{T}^{-1}b(c).$ (H.1)
Denote
$a_{\ell}:=S_{\mathcal{A}}W^{\ell}e_{j}\in\mathbb{R}^{m},\qquad\ell\in\mathbb{N}.$
Then for $t=1,\dots,T$, the $t$th $m$-block of $b(c)$ satisfies the same convolution identity
$b_{t}(c)=\sum_{k=1}^{t}a_{t-k}\,c_{k}.$ (H.2)
We start by decomposing $K_{T}$. By Lemma 9 (Lemma 9 in Appendix C), we can write
$W=P+R,\qquad P:=\frac{1}{n}\mathbf{1}\mathbf{1}^{\top},\qquad\mathbf{1}^{\top}R=0,\qquad RP=PR=0,$ (H.3)
and $\|R^{t}\|_{2}\leq\rho^{t}$ for $\rho:=\max_{k\geq 2}|\lambda_{k}(W)|\in(0,1)$. Using (H.3) together with the block-Toeplitz structure of $H_{T}$ and the fact that $S_{\mathcal{A}}S_{\mathcal{A}}^{\top}=I_{m}$, one obtains the decomposition
$K_{T}=\frac{1}{n}(J\otimes I_{m})+G_{T},$ (H.4)
where $G_{T}\succeq 0$ and $J\succ 0$ satisfies $J=LL^{\top}$ with $L$ the $T\times T$ lower-triangular all-ones matrix. Equivalently, $J_{t,s}=\min(t,s)$. Therefore,
$K_{T}\succeq\frac{1}{n}(J\otimes I_{m})$
which further implies
$K_{T}^{-1}\preceq n(J^{-1}\otimes I_{m}).$ (H.5)
Let $D$ be the backward difference matrix on $\mathbb{R}^{T}$,
$(Dx)_{1}=x_{1},\qquad(Dx)_{t}=x_{t}-x_{t-1}\ \ (t\geq 2),$
so that $J^{-1}=D^{\top}D$ and for all $x\in\mathbb{R}^{T}$,
$x^{\top}J^{-1}x=\|Dx\|_{2}^{2}.$ (H.6)
We next split $b(c)$ into stationary and transient parts, exactly as in the proof of Thm. 13. Define the stationary and transient terms as
$a_{\ell}^{(P)}:=S_{\mathcal{A}}Pe_{j}=\frac{1}{n}\mathbf{1}_{m}\quad(\ell\geq 1),\qquad a_{\ell}^{(R)}:=S_{\mathcal{A}}R^{\ell}e_{j}\quad(\ell\geq 1),$ (H.7)
and $a_{0}=S_{\mathcal{A}}e_{j}$. Accordingly, write
$b(c)=b^{(0)}(c)+b^{(P)}(c)+b^{(R)}(c),$ (H.8)
for $t=1,\dots,T$,
$b_{t}^{(P)}(c):=\sum_{k=1}^{t-1}\frac{1}{n}\,\mathbf{1}_{m}\,c_{k}=\frac{1}{n}\mathbf{1}_{m}\sum_{k=1}^{t-1}c_{k},$ (H.9)
and
$b_{t}^{(R)}(c):=\sum_{m=1}^{t-1}S_{\mathcal{A}}R^{m}e_{j}\,c_{t-m}.$ (H.10)
We next derive an upper bound for the quadratic form $b(c)^{\top}K_{T}^{-1}b(c)$. For any $c$,
$b(c)^{\top}K_{T}^{-1}b(c)$ $=\|b^{(0)}(c)+b^{(P)}(c)+b^{(R)}(c)\|_{K_{T}^{-1}}^{2}$
$\leq\Big(\|b^{(0)}(c)\|_{K_{T}^{-1}}+\|b^{(P)}(c)\|_{K_{T}^{-1}}+\|b^{(R)}(c)\|_{K_{T}^{-1}}\Big)^{2}.$
By (H.5) and (H.6),
$\|b^{(P)}(c)\|_{K_{T}^{-1}}^{2}$ $\leq n\,(b^{(P)}(c))^{\top}(J^{-1}\otimes I_{m})\,b^{(P)}(c)=n\,\|(D\otimes I_{m})b^{(P)}(c)\|_{2}^{2}.$
From (H.9), we have $(Db^{(P)}(c))_{t}=\frac{1}{n}\,\mathbf{1}_{m}\,c_{t-1}$ for $t\geq 2$ and $(Db^{(P)}(c))_{1}=0$, hence
$\|(D\otimes I_{m})b^{(P)}(c)\|_{2}^{2}=\sum_{t=2}^{T}\left\|\frac{1}{n}\,\mathbf{1}_{m}\,c_{t-1}\right\|_{2}^{2}=\sum_{t=2}^{T}\frac{m}{n^{2}}=\frac{m(T-1)}{n^{2}}.$
Therefore
$\|b^{(P)}(c)\|_{K_{T}^{-1}}^{2}\leq\frac{m(T-1)}{n}.$ (H.11)
By Lemma H.1, $\|b^{(R)}(c)\|_{K_{T}^{-1}}^{2}\leq C_{\mathrm{mix}}(W)$. Thus, for all $c$,
$b(c)^{\top}K_{T}^{-1}b(c)\leq\left(\sqrt{\frac{m(T-1)}{n}}+\sqrt{C_{\mathrm{mix}}(W)}+O(1)\right)^{2}=\frac{m}{n}T+O(\sqrt{T}).$
The constants given in the theorem statement are obtained by combining the $O(1)$ term with the $\sqrt{\frac{m(T-1)}{n}}$-term. ∎
H.2 Bound for the Transient Term
We next prove the required technical lemma which shows that the transient terms stay bounded independent of $T$ also for the multi-node observation.
Lemma H.1.
Let $c\in\{-1,1\}^{T}$ and let $\widetilde{b}^{(R)}(c)$ and $\widetilde{K}_{T}$ be defined as above in the proof of Thm. 15. There exists a constant $C_{\mathrm{mix}}(W)<\infty$ depending only on $W$, such that for all $T\geq 1$ and all $c\in\{\pm 1\}^{T}$,
$\left(\widetilde{b}^{(R)}(c)\right)^{\top}\widetilde{K}_{T}^{-1}\,\widetilde{b}^{(R)}(c)\leq C_{\mathrm{mix}}(W),$ (H.12)
where one may take
$C_{\mathrm{mix}}(W)=\frac{\pi^{2}}{6}\,m\,(1-\rho)^{-4},\qquad\rho:=\max_{k\geq 2}|\lambda_{k}(W)|.$
Proof.
The proof follows the same Schur-complement argument as Lemma G.2, with the only difference that now the vectors live in $\mathbb{R}^{mT}$.
For $t=1,\dots,T$, let $\widetilde{K}_{t}$ be the leading $(mt)\times(mt)$ principal submatrix of $\widetilde{K}_{T}$ and write it in block form
$\widetilde{K}_{t}=\begin{pmatrix}\widetilde{K}_{t-1}&\widetilde{k}_{t}\\ \widetilde{k}_{t}^{\top}&\widetilde{\kappa}_{t}\end{pmatrix},\qquad\widetilde{s}_{t}:=\widetilde{\kappa}_{t}-\widetilde{k}_{t}^{\top}\widetilde{K}_{t-1}^{-1}\widetilde{k}_{t}\succ 0,$
where $\widetilde{\kappa}_{t}\in\mathbb{R}^{m\times m}$ and $\widetilde{s}_{t}\in\mathbb{R}^{m\times m}$. For any $x\in\mathbb{R}^{mT}$ define the residual vector
$\widetilde{r}_{t}(x):=x_{t}-\widetilde{k}_{t}^{\top}\widetilde{K}_{t-1}^{-1}x_{1:t-1}\in\mathbb{R}^{m},$
where $x_{t}\in\mathbb{R}^{m}$ denotes the $t$th block of $x$. A standard block inverse identity with Schur complement gives
$x^{\top}\widetilde{K}_{T}^{-1}x=\sum_{t=1}^{T}\widetilde{r}_{t}(x)^{\top}\widetilde{s}_{t}^{-1}\widetilde{r}_{t}(x).$ (H.13)
We next lower bound $\widetilde{s}_{t}$. Let $\widetilde{h}_{t}^{\top}$ be the $t$th block-row of $\widetilde{H}_{T}$. Then, $\widetilde{K}_{T}$ is the Gram matrix of $\{\widetilde{h}_{t}\}_{t=1}^{T}$. Here $\widetilde{s}_{t}$ admits a geometric interpretation analogous to the scalar case. Since $\widetilde{K}_{T}$ is the Gram matrix of the block-rows $\{\widetilde{h}_{1},\dots,\widetilde{h}_{T}\}$ of $\widetilde{H}_{T}$, the Schur complement
$\widetilde{s}_{t}=\widetilde{\kappa}_{t}-\widetilde{k}_{t}^{\top}\widetilde{K}_{t-1}^{-1}\widetilde{k}_{t}$
is the Gram matrix of the component of $\widetilde{h}_{t}$ orthogonal to $\mathrm{span}\{\widetilde{h}_{1},\dots,\widetilde{h}_{t-1}\}$. Equivalently, for any $z\in\mathbb{R}^{m}$, $z^{\top}\widetilde{s}_{t}z$ equals the squared distance of the scalar row $z^{\top}\widetilde{h}_{t}$ from the span of $\{z^{\top}\widetilde{h}_{1},\dots,z^{\top}\widetilde{h}_{t-1}\}$, and $\widetilde{s}_{t}$ equals the squared-distance operator (in $\mathbb{R}^{m}$) from $\widetilde{h}_{t}$ to $\mathrm{span}\{\widetilde{h}_{1},\dots,\widetilde{h}_{t-1}\}$. Since $\widetilde{H}_{T}$ is block lower-triangular with diagonal block $US_{\mathcal{A}}$, and $US_{\mathcal{A}}$ has orthonormal rows (because $S_{\mathcal{A}}S_{\mathcal{A}}^{\top}=I_{m}$ and $U$ is orthogonal), the diagonal block is orthogonal to the past block-rows, implying
$\widetilde{s}_{t}\succeq I_{m}\quad\Rightarrow\quad\widetilde{s}_{t}^{-1}\preceq I_{m}.$ (H.14)
Combining (H.13)–(H.14) gives
$x^{\top}\widetilde{K}_{T}^{-1}x\leq\sum_{t=1}^{T}\|\widetilde{r}_{t}(x)\|_{2}^{2}.$ (H.15)
We now bound the residuals for $x=\widetilde{b}^{(R)}(c)$. As in the proof of Lemma G.2, the residual $\widetilde{r}_{t}(x)$ is the error of the best linear predictor of $x_{t}$ from the past $x_{1:t-1}$. Hence for any (possibly suboptimal) predictor $\widehat{x}_{t}$ measurable w.r.t. the past,
$\|\widetilde{r}_{t}(x)\|_{2}^{2}\leq\|x_{t}-\widehat{x}_{t}\|_{2}^{2}.$ (H.16)
Fix $t\geq 2$ and set the truncation length
$m_{t}:=\left\lceil\frac{\log t}{\log(1/\rho)}\right\rceil\quad\Rightarrow\quad\rho^{m_{t}}\leq\frac{1}{t}.$
We use the same predictor as in the proof of Lemma G.2, applied component-wise:
$\widehat{\widetilde{b}}^{(R)}_{t}:=\sum_{q=1}^{m_{t}}\beta_{q}\,\widetilde{b}^{(R)}_{t-q}(c),\qquad\beta_{q}:=(W^{q})_{ii},$ (H.17)
for an arbitrary but fixed $i\in\mathcal{A}$ (any $i$ works; only $|\beta_{q}|\leq 1$ is used).
Using same techniques as in the proof of Lemma G.2, i.e., using (i) $W^{q}R^{\ell-q}=R^{\ell}$, (ii) row-stochasticity, and (iii) $\|R^{k}\|_{2}\leq\rho^{k}$, gives the uniform tail bound
$\left\|\widetilde{b}^{(R)}_{t}(c)-\widehat{\widetilde{b}}^{(R)}_{t}\right\|_{2}\leq\frac{\sqrt{m}}{(1-\rho)^{2}}\,\frac{1}{t}.$ (H.18)
Here the extra factor $\sqrt{m}$ comes from bounding an $\ell_{2}$-norm over $m$ coordinates by $\sqrt{m}$ times a coordinate-wise bound.
Combining (H.16) and (H.18), we obtain
$\|\widetilde{r}_{t}(\widetilde{b}^{(R)}(c))\|_{2}^{2}\leq\left\|\widetilde{b}^{(R)}_{t}(c)-\widehat{\widetilde{b}}^{(R)}_{t}\right\|_{2}^{2}\leq\frac{m}{(1-\rho)^{4}}\,\frac{1}{t^{2}}.$ (H.19)
Using (H.15), summing over $t$ and using the bound (H.19) gives
$\left(\widetilde{b}^{(R)}(c)\right)^{\top}\widetilde{K}_{T}^{-1}\,\widetilde{b}^{(R)}(c)\leq\sum_{t=1}^{T}\|\widetilde{r}_{t}(\widetilde{b}^{(R)}(c))\|_{2}^{2}\leq\frac{m}{(1-\rho)^{4}}\sum_{t=1}^{\infty}\frac{1}{t^{2}}\leq\frac{\pi^{2}/6}{(1-\rho)^{4}}\,m.$
This proves the claim. ∎
Appendix I Extension to the Case When Observers’ Noise Does not Contribute (SecLDP)
We next extend the proof technique of Theorem 15 to the SecLDP setting, where the adversary observing a set $\mathcal{A}$ can remove the Gaussian noise injected by nodes in $\mathcal{A}$. Equivalently, one deletes the corresponding columns in each time-block column of the system matrix.
I.1 Proof of Theorem 16
We simply sketch the proof as it has exactly the same structure as the proof of Theorem 15.
In the SecLDP setting, the adversary observing $\mathcal{A}$ can remove the Gaussian noises injected by nodes in $\mathcal{A}$. This is modeled by right-multiplying each time-block column of the system matrix by
$P_{\mathcal{A}^{c}}:=I_{n}-\sum_{i\in\mathcal{A}}e_{i}e_{i}^{\top},\qquad\bar{H}_{T}:=H_{T}(I_{T}\otimes P_{\mathcal{A}^{c}}),\qquad\bar{K}_{T}:=\bar{H}_{T}\bar{H}_{T}^{\top}.$
For $j\notin\mathcal{A}$, $(I_{T}\otimes P_{\mathcal{A}^{c}})(c\otimes e_{j})=c\otimes e_{j}$, hence: $\bar{b}(c):=\bar{H}_{T}(c\otimes e_{j})=H_{T}(c\otimes e_{j})=b(c)$. Therefore the SecLDP sensitivity is
$\bar{\Delta}_{T}(j\to\mathcal{A})^{2}=\max_{c\in\{\pm 1\}^{T}}b(c)^{\top}\bar{K}_{T}^{+}\,b(c),$
and the difference to Theorem 15 is in the modified Gram matrix $\bar{K}_{T}$.
We decompose $W=P+R$ as in the single- and multi-node proofs, with $P=\frac{1}{n}\mathbf{1}\mathbf{1}^{\top}$, $PR=RP=0$, and $\|R^{t}\|_{2}\leq\rho^{t}$. This gives a decomposition $\bar{H}_{T}=\bar{H}_{T}^{(P)}+\bar{H}_{T}^{(R)}$ and, expanding,
$\bar{K}_{T}=\bar{K}_{T}^{(PP)}+\bar{K}_{T}^{(RR)}+\bar{K}_{T}^{(PR)}+\bar{K}_{T}^{(RP)},\qquad\bar{K}_{T}^{(PP)}:=\bar{H}_{T}^{(P)}(\bar{H}_{T}^{(P)})^{\top},\ \ \bar{K}_{T}^{(RR)}:=\bar{H}_{T}^{(R)}(\bar{H}_{T}^{(R)})^{\top}.$
Unlike in the proof of Theorem 15, the cross terms $\bar{K}_{T}^{(PR)}+\bar{K}_{T}^{(RP)}$ do not vanish in general because of the right-multiplication by $P_{\mathcal{A}^{c}}$ (we cannot use the orthogonality $\mathbf{1}^{\top}R=0$). We keep these terms explicitly and control them by Cauchy–Schwarz: for any $x$ and any $\eta>0$,
$\big|x^{\top}(\bar{K}_{T}^{(PR)}+\bar{K}_{T}^{(RP)})x\big|\leq\eta\,x^{\top}\bar{K}_{T}^{(PP)}x+\eta^{-1}x^{\top}\bar{K}_{T}^{(RR)}x,$
so the cross terms can be absorbed into the stationary and transient terms.
The stationary term is computed explicitly. Using $P=\frac{1}{n}\mathbf{1}\mathbf{1}^{\top}$ and $P_{\mathcal{A}^{c}}\mathbf{1}=\mathbf{1}_{\mathcal{A}^{c}}$, one obtains the Toeplitz form
$\bar{K}_{T}^{(PP)}=\frac{n-m}{n^{2}}\,(J\otimes\mathbf{1}_{m}\mathbf{1}_{m}^{\top}),\qquad J_{t,s}=\min(t,s),$
i.e., compared to the proof of Theorem 15, we get the coefficient (nm)/n(n-m)/n because only the $\mathcal{A}^{c}$ noise coordinates remain. Consequently, the same backward-difference argument ($J^{-1}=D^{\top}D$) used in (H.11) gives the SecLDP linear term with slope $\frac{m}{n-m}$ (instead of $\frac{m}{n}$).
Finally, all terms involving $R$ (the pure transient contribution and the cross-term contribution) are controlled by the exactly the same Schur-complement technique as Lemma G.2 and Lemma H.1, giving a uniform bound independent of $T$:
$(b^{(R)}(c))^{\top}\bar{K}_{T}^{+}b^{(R)}(c)\leq C_{\mathrm{mix}}(W),$
and similarly for the cross terms. Putting the stationary and transient terms together gives the claim.
Appendix J Proofs for Section 5 (Utility Analysis)
J.1 From Absolute Spectral Gap to the Consensus Parameter $p$
We prove the standard contraction inequality and its implication for the parameter $p$ that is needed in the unified decentralized SGD analysis of (Koloskova et al., 2020).
Lemma J.1.
Let $W\in\mathbb{R}^{n\times n}$ be symmetric and doubly-stochastic. Let $1=\lambda_{1}(W)\geq\lambda_{2}(W)\geq\cdots\geq\lambda_{n}(W)\geq-1$ be its eigenvalues. For any $X\in\mathbb{R}^{d\times n}$, define $\bar{X}:=X\frac{1}{n}\mathbf{1}\mathbf{1}^{\top}$ and $\rho=\max\{|\lambda_{2}(W)|,|\lambda_{n}(W)|\}$. Then
$\|XW-\bar{X}\|_{F}^{2}\;\leq\;\rho^{2}\,\|X-\bar{X}\|_{F}^{2}.$ (J.1)
Consequently, the assumption
$\|XW-\bar{X}\|_{F}^{2}\leq(1-p)\|X-\bar{X}\|_{F}^{2}$ (J.2)
needed in the analysis of (Koloskova et al., 2020) holds with $p=1-\rho^{2}$.
Proof.
Let $P:=\frac{1}{n}\mathbf{1}\mathbf{1}^{\top}$ be the orthogonal projector onto $\mathrm{span}\{\mathbf{1}\}$. Then $\bar{X}=XP$ and, since $W\mathbf{1}=\mathbf{1}$ and $\mathbf{1}^{\top}W=\mathbf{1}^{\top}$, we have $WP=PW=P$. Hence
$XW-\bar{X}=XW-XP=X(W-P)=(X-\bar{X})(W-P),$
because $\bar{X}(W-P)=XP(W-P)=X(PW-P)=0$. Now, WPW-P is symmetric with eigenvalues 0 (along $\mathbf{1}$) and $\lambda_{2}(W),\ldots,\lambda_{n}(W)$ on $\mathbf{1}^{\perp}$, so $\|W-P\|_{2}=\max\{|\lambda_{2}(W)|,|\lambda_{n}(W)|\}=\rho$ (since $W$ is symmetric). Thus
$\|XW-\bar{X}\|_{F}=\|(X-\bar{X})(W-P)\|_{F}\leq\|X-\bar{X}\|_{F}\,\|W-P\|_{2}=\rho\,\|X-\bar{X}\|_{F}.$
Using the relation $1-\rho^{2}=(1+\rho)(1-\rho)$ we directly get the following auxiliary result.
Lemma J.2.
Let $W$ be symmetric and doubly-stochastic and define the absolute spectral gap
$\gamma:=1-\max\{|\lambda_{2}(W)|,|\lambda_{n}(W)|\}\in(0,1].$
With $p=1-\rho^{2}$,
$\frac{1}{p}=O\!\left(\frac{1}{\gamma}\right),\;\;\frac{1}{p^{2}}=O\!\left(\frac{1}{\gamma^{2}}\right).$ (J.3)
J.2 Deriving the Utility Bound from Results of (Koloskova et al., 2020)
We derive Theorem 17 by plugging our privacy calibration into a strongly-convex convergence bound by (Koloskova et al., 2020). More specifically, we use the following template bound (a specialization of the unified analysis to $\tau=1$ and strongly convex objectives).
Lemma J.3 (Koloskova et al. 2020).
[…]